OpenStudy (lucaz):

function f, f(1) = 2, f(x+1) = f(x)-1, f(100)=?

OpenStudy (amistre64):

isnt this just recurrsion?

OpenStudy (amistre64):

\[a_n=a_{n-1}-1\]

OpenStudy (amistre64):

2 - 1(0) 2 - 1(1) 2 - 1(2) 2 - 1(3) .... 2 - 1(100)

OpenStudy (amistre64):

may 2-1(99)

OpenStudy (lucaz):

I don't get how it works

OpenStudy (amistre64):

its like working in an office, youjust do the same thing over and over and over again

OpenStudy (amistre64):

f(1) = 1 f(x+1) = f(x) - 1 f(1+1) = f(1) - 1; f(2) = 2-1 = 1 f(2+1) = f(2) - 1 = 1 - 1 = 0 etc ....

OpenStudy (amistre64):

well, f(1) = 2 .... wrong key lol

OpenStudy (lucaz):

f(1)=2, not 1

OpenStudy (amistre64):

you can either do that 100 times, or develop the pattern: f(x) = 2 - (x-1)

OpenStudy (lucaz):

why f(x)=2-(x-1)? sorry..

OpenStudy (amistre64):

its a pattern that develops over the course of the repetitions .. f(1) = 2 f(2) = f(1) - 1= 2 - 1 f(3) = f(2) - 1= f(1) -1 - 1 = 2 - 1 - 1 f(4) = f(3) - 1= f(1) -1 - 1 - 1 = 2 - 1 - 1 - 1 .... f(n) = 2 - (1 - 1 - 1 - ... - 1), n-1 times f(n) = 2 - (n-1)

OpenStudy (lucaz):

right, I get it, thanks!

OpenStudy (amistre64):

youre welcome

OpenStudy (lucaz):

can you help with other AP question?

OpenStudy (amistre64):

maybe one more

OpenStudy (lucaz):

guy A walks 10km/day, the guy B walks 8km the first day and adds 0.5 each day to reach the guy A, how many days it took?

OpenStudy (amistre64):

i assume B is trying to catch upto A?

OpenStudy (lucaz):

yes

OpenStudy (amistre64):

lets assume the start at the same place, when does their distances equal out? d1: 10: 8 d2: 20: 8 + 8 + .5 d3: 30: 8 + 8 + .5 + 8 + .5 + .5 d4: 40: 8 + 8 + .5 + 8 + .5 + .5 + 8 + .5 + .5 + .5 or written another way d1: 10: 8 d2: 2(10): 2(8) + .5 d3: 3(10): 3(8) + .5 + .5 + .5 d4: 4(10): 4(8) + .5 + .5 + .5 + .5 + .5 + .5 dn: n(10) = n(8) + 1(.5) + 2(.5) + 3(.5)+ ... + (n-1)(.5) 2n = .5(1+2+3+...+(n-1)) 4n = 1+2+3+...+(n-1) \[4n = \frac{(n-1)(n)}{2}\]\[4 = \frac{n-1}{2}\]\[8 = n-1\]\[9 = n\] so my guess is in 9 days

OpenStudy (amistre64):

after 8 days, that 80 to 79, so on the 9th day they catch up

OpenStudy (lucaz):

where did that 2n = .5(1+2+3+...+(n-1)) come from?

OpenStudy (lucaz):

0.5 is the common factor..

OpenStudy (amistre64):

from the information given, each day the speed is .5 faster than the day before day : distance covered 1 8 2 8 + .5 3 8 + .5 + .5 4 8 + .5 + .5 + .5 the distance covered over 4 days is the sum of all the days 4(8) + 0(.5) + 1(.5) + 2(.5) + 3(.4) 4(8) + .5 (1 + 2 + 3) on the nth day 8n + .5 (1 + 2 + 3 + ... + n-1)

OpenStudy (amistre64):

since the distance covered by each guy has to be equal for them to be at the same place; the other guy is walking 10 each day; 10n 10n = 8n + .5(....) 2n = .5(....) 4n = (....)

OpenStudy (lucaz):

I see

OpenStudy (lucaz):

then it's just algebra, ok, thanks again!

OpenStudy (amistre64):

its always just algebra ;) good luck

OpenStudy (lucaz):

=]